OpenStudy (anonymous):

]Choose the equation of the line passing through the point (-1, -2) and perpendicular to y =1/4 x + 2. y = 4x - 6 y = -4x - 6 y = -4x + 2 y = 4x

OpenStudy (anonymous):

in order to be perpendicular than you invert m and multiply by -1 by m I mean the m in y=mx+b your equation is already in this mode so m is 1/4 invert it and multiply by -1, then it is - 4/1

OpenStudy (anonymous):

so now if you want it to pass thru -1,-2 plug it into y=mx+b -2=-4/1*-1+b solve for b and b =-6 therfore y=-4/1-6

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

how bout this. Which of the following is the slope between the points (0, -10) and (-4, 0)? 2/5 -2/5 -5/2 5/2

OpenStudy (anonymous):

slope is (y2-y1)/(x2-x1)=(0+10)/(-4)=-10/4 -5/2

OpenStudy (anonymous):

thanks how about this . ]Which of the following is the slope between the points (-1, -5) and (7, -3)? 4 1/4 -1/4 -4

OpenStudy (anonymous):

1/4 is the slope

OpenStudy (anonymous):

slope is (y2-y1)/(x2-x1)=(-3-(-5))/(7-(-1))

OpenStudy (anonymous):

the formula used by timo86m gives the solution

OpenStudy (anonymous):

slope is (y2-y1)/(x2-x1) most times it is written slope is (yf-yi)/(xf-xi) the f and i means final and initial.

OpenStudy (anonymous):

its easier to remember that final- initial